### v1

@@ 1 Introduction
The rapid development and deployment of AI systems has made questions of safety increasingly urgent, demanding immediate attention from policymakers and governance bodies as they face critical decisions about regulatory frameworks, liability standards, and safety certification requirements for AI systems. Recent discourse has narrowed to focus primarily on AI safety as a project of minimizing existential risks from future advanced AI systems [1, 2, 3]. [^1] The focus on existential risks from AI has been particularly connected to normative theories and movements such as rationalism, effective altruism, and longtermism [4, 5, 6]. While these theories offer valuable perspectives on long-term challenges, their specific institutional articulations have faced substantial criticism [7, 8, 9]. As the relationship between AI safety and these normative theories and movements is quite complex, the association between AI safety and the conventional conception of AI existential risk is commonly presented as a defining characteristic [10, 11, 12, 3, 2]. This situation is particularly concerning to researchers and practitioners who work on the development and deployment of safe and responsible AI, but do not align with these normative theories or existential risk narratives posed by AI [13, 14, 15, 16]. This disparity has led to a divisive and sometimes unhealthy atmosphere, with some researchers even questioning the unique contribution of “AI safety” community [17]. This concentrated attention on existential risk has emerged despite — and perhaps overshadowed — decades of engineering and technical progress in building robust and reliable AI systems.
@@ 6 Discussion and future research
In this paper, we conducted a systematic review of primarily peer-reviewed literature to unpack the diverse technical and practical challenges encompassed by the term “safe AI.” Our empirical analysis shows a broad landscape of motivations and outcomes driving AI safety research. The significance of these motivations and research outcomes stems from a desire to ensure that the AI systems we are building are reliable, trustworthy, and beneficial for society. By examining a diverse body of peer-reviewed literature, we have found that AI safety research addresses a wide range of risks across the entire lifecycle of AI systems.
@@
Several research questions remain open. We highlight three of them below.
First, the concept of “sociotechnical AI safety” and its implications for the field of AI safety further investigation. While recent work [179] has begun to clarify this notion in the context of AI evaluation practices, more research is needed to examine the ramifications of sociotechnical approaches for AI safety research and practice as a whole. This includes exploring how to effectively integrate social, ethical, and political considerations into technical safety research and development.

### v2

@@ 1 Introduction
The rapid development and deployment of AI systems has made questions of safety increasingly urgent, demanding immediate attention from policymakers and governance bodies as they face critical decisions about regulatory frameworks, liability standards, and safety certification requirements for AI systems. Recent discourse has narrowed to focus primarily on AI safety as a project of minimizing existential risks from future advanced AI systems [1, 2, 3]. [^1] The focus on existential risks from AI has been particularly connected to normative theories and movements such as rationalism, effective altruism, or longtermism [5, 6]. While these theories offer valuable perspectives on long-term challenges, their specific institutional articulations have faced substantial criticism [7, 8, 9]. As the relationship between AI safety and these normative theories and movements is quite complex, the association between AI safety and the conventional conception of AI existential risk is commonly presented as a defining characteristic [10, 11, 12, 3, 2]. This situation is particularly concerning to researchers and practitioners who work on the development and deployment of safe and responsible AI, but do not align with these normative theories or existential risk narratives posed by AI [13, 14, 15, 16]. This disparity has led to a divisive and sometimes unhealthy atmosphere, with some researchers even questioning the unique contribution of “AI safety” community [17]. This concentrated attention on existential risk has emerged despite — and perhaps overshadowed — decades of engineering and technical progress in building robust and reliable AI systems.
@@ 6 Discussion and future research
In this paper, we conducted a systematic review of primarily peer-reviewed literature to unpack the diverse technical and practical challenges encompassed by AI safety. Our empirical analysis shows a broad landscape of motivations and outcomes driving AI safety research. The significance of these motivations and research outcomes stems from a desire to ensure that the AI systems we are building are reliable, trustworthy, and beneficial for society. By examining a diverse body of peer-reviewed literature, we have found that AI safety research addresses a wide range of risks across the entire lifecycle of AI systems.
@@
Several research questions remain open. We highlight two of them below.
First, the concept of “sociotechnical AI safety” and its implications for the field of AI safety require further investigation [2]. Although recent work [179] has begun to clarify this notion in the context of AI evaluation practices, more research is needed to examine the implications of sociotechnical approaches for AI safety research and practice as a whole. This includes exploring how to effectively integrate social, ethical, and political considerations into technical safety research and development.